A dishonest milk vendor purchased pure milk at Rs.10 per liter and sold diluted milk at Rs.10 per liter and gains a profit of 60%. What quantity of pure milk did he use in every liter?
Answer by ghostrecog(4)

Let the vendor purchased a 10 liters of milk at Rs.10 per liter.
Then the total amount of purchase is 100.
he gained 60% profit after diluting.
Let x be rate of milk per liter.
and y be the total amount of liters
and w be the water added in order to dilute.
%gain = (new amount-old amount)/old amount
0.6= (x(y+w)-xy)/xy
w/y=0.6
w/y=6/10=3/5
This is clearly states that out of 8 liters, 3 liters of water is added with 5 liters of milk.
As per above statement total diluted milk =16 liters.
So there is 10 liters of pure milk in 16 liters of diluted milk.
Therefore each liter has a pure milk of =10/16 liters
=0.625 liters
=625 ml